A money lender borrows money at 4% pa and paid the interest at the end of the year. He lent it at 6% pa CI compounded half yearly. In this way he gains 104.50Rs. Find the amount of money he borrowed?

Grade:8

2 Answers

Seema
8 years ago
Amount= p [1+r/100]^no of years.
First case: 
A1=P [1+4/100]^1
Second case:
Since its compounded half yearly so time is doubled and rate is halved.
A2=P [1+3/100]^2
So,
104.50=A2-A1. 
Solve the equation to get the answer
Nivi
5 years ago
Let the borrow amount be =X
Given,X [(1+3/100)2-1]-X×4×1/100=104.50
 
Interest compounded half yearly
R=6/2=3  ,  N=1×2=2
 
X [(1.03)2-1]-00.00404X =104.50
 
0.0609X -0.04X =104.50
 
0.0209X =104.50
X = 104.50/0.0209
X =5000